OpenStudy (anonymous):

At the local convenience store, 3 bags of chips and 2 containers of dip cost $9. However, 5 bags of chips and 7 containers of dip cost $26. What is the cost of one bag of chips and one container of dip?

OpenStudy (anonymous):

make two eqns

OpenStudy (anonymous):

3x+2y=59 5x+7y=26

OpenStudy (anonymous):

then solve:: do u need further help..

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

did u get the two eqns'

OpenStudy (anonymous):

from eqn i> x=(59-2y)/3 substitute this value of x in second eqn 5x+7y=26 5 (59-2y)/3 +7y=26 (295-10y)/3+7y=26 295-10y+21y=26 X 3 11y=78-295 y=-217/11 x=(59-2y)/3 =(59-2 X(-217/11))/3 =(649-6293)/33 y=-5644/33

OpenStudy (anonymous):

sry x=-5644/33

OpenStudy (anonymous):

there may be some cal errors

OpenStudy (anonymous):

Woha! that's alot! a. chips = $2; dip = $2 b. chips = $1.50; dip = $2.50 c. chips = $3; dip = $4 d. chips = $1; dip = $3 these are the choices and i got b

OpenStudy (anonymous):

x is the cost of one bag of chips and y is the cost of one bag of dip

OpenStudy (anonymous):

in the first eqn it ought to be 9 instead of 59..

OpenStudy (anonymous):

so just substitute it....................hopefully ul get the answer . . . there is also an alternative method using matrix.. but that difficult to show

OpenStudy (anonymous):

ans is d>

OpenStudy (anonymous):

i'v don it finally..

OpenStudy (anonymous):

do you need the process......

OpenStudy (anonymous):

u got confused ??

OpenStudy (anonymous):

cost of a bag of chip is 1 and of a box of dip is 3

OpenStudy (anonymous):

Just a bit but i think i got it know! thank you!!!

OpenStudy (anonymous):

ur welllcommm..